OpenStudy (anonymous):

Evaluate each expression for a = -2 and b = 6. 1.2a-1b-2 2.-4a-2 b-3 3.3-2a-2b-1

OpenStudy (hoblos):

just replace a & b by their values i'll solve the first one and you try the others 1) 2(-2) - 6 -2 = -4-6-2=-12

OpenStudy (anonymous):

from where you got the -6?

OpenStudy (hoblos):

-1b=-1(6)=-6

OpenStudy (anonymous):

ill solve 2nd one check it plz. -4(-2)-2(6)-3=-8-12-3=-23

OpenStudy (hoblos):

-4(-2)= +8 !!

OpenStudy (anonymous):

=-7